Dit is de Titel

M. van Klonkerdom April 18, 2002

Abstract In dit voorbeeld wordt een aantal elementen uit LaTeX getoond: graphics, vergelijkin-

gen, en bibliografische referenties komen aan de orde.

1 Introduction

Bla bla bla zoals onderzocht in [2]. Output (action) is determined by

A_{t} = F(A_{[t_{0},t−δt]}, P_{[t_{0},t]})  (1)

where P_{[t_{0},t]} is the perceptual history, A_{[t_{0},t−δt]} is the action history of the agent which came into existence at t_{0}...

2 Method

Bla bla in Figuur 1 we see ....

Figure 1: Een schema

3 Results

Bla bla. The concept of metastable states [1] has been thorougly ...

4 Conclusion

It was ok.

References

[1] E. Gardner, Structure of metastable states in the Hopfield model J. Phys. A: Math. Gen. 19 (1986) L1047

[2] D.H.Ackley, G.E.Hinton, T.J. Seinowski Cognitive Sci. 9 (1985) 147